Yoshi is 12 years older than his sister. Six years from now, the sum of their age will be 32. Find Yoshi's present age.
Found 2 solutions by mananth, Theo:
Answer by mananth(16946)

sister = x years
yoshi = x+12
..
after six years
sister=x+6
Yoshi = x+12+6=x+18
..
x+18+x+6 = 32
2x+24=32
-24 -24
2x=32-24
2x=8
/2
x=4 sister's age
Yoshi's age = 12+4 = 16 years
...
m.ananth@hotmail.ca

Answer by Theo(13342)

Let x = yoshi's sister's age.
yoshi's age is x + 12.

6 years from now the sum of their age will be 32.

6 years from now yoshi's sister will be x + 6 years old.

6 years from now yoshi will be (x + 12) + 6 years old.

The sum of their ages will be x + 6 + (x + 12) + 6 = 2*x + 24 = 32

You have 2*x + 24 = 32

Solve for x to get x = 4

yoshi's sister's age today is 4.
in 6 years yoshi's sister will be 10 years old.

yoshi is 16 years old today.
in 6 years, yoshi will be 22 years old.

10 + 22 = 32

answer to the question is that yoshi's present age is 16 years old.